# Position Paper Are Normalizing Flows the Key to Unlocking the Exponential Mechanism? A Path through the Accuracy-Privacy Ceiling Constraining Differentially Private ML

### v3

@@ 2.2 Normalizing Flows (NFs)
  Initialize $\theta_{\text{NF},0}$ randomly
  for $t$ in $[T]$Train Auxilary NF Model} do
     Generate $N$ samples from base distribution
     $z_{i}\sim p_{z}$
     Produce $N$ parameter samples & Jacobian
     $\theta_{i}=g_{\theta_{\text{NF},t}}(z_{i})$
@@
     Compute utility for each parameter sample
@@
     Compute Reverse KL Loss using Eq. 1
@@
     $\theta_{\text{NF},t+1}=\theta_{\text{NF},t}-\eta_{t}\nabla_{\theta_{\text{NF},t}}\mathcal{L}$
@@
  Output Model parameters, $\theta=g_{\theta_{\text{NF},T}}(z)$, privacy $\varepsilon$
3 ExpM+NF Method
Figure 1 describes the setting that we will consider: a private data enclave that wishes to permit optimization on their data; in particular, training and publicly releasing an ML algorithm on that data. The basic method for producing a differentially private ML model with ExpM+NF is given in Algorithm 1.
Fitting an ML model on training features and targets $(X,Y)$ constitutes finding parameters $\theta$ that minimizes a loss function $L(X,Y,\theta)$, ExpM requires specifying a utility function $u(X,Y,\theta)$ to be maximized in $\theta$; thus, $u=\phi\circ L$ for strictly decreasing $\phi$. Throughout, we will consider $u=-L$111 While any strictly decreasing $\phi$ theoretically ensures identical optimization, in practice $\phi$ must separate the near-optimal $\theta$ from those that are far from optimal. This means the derivative must be a sufficiently large negative value over an appropriate range of $L(\theta)$ to discriminate high versus low utility $\theta$. In particular, using a negative sigmoid function, which is flat ($\phi^{\prime}\approxeq 0$) except in a limited neighborhood, is unwise, unless we know which range of $L(\theta)$ will contain the optimal values a priori and shift $\phi$ to discriminate these values. Lastly, we note that choice of $\phi$ effects the sensitivity, and therefore may be designed to assist in producing the required sensitivity bound. To define the ExpM density, we must provide $s$, a bound on $u$’s sensitivity (Defn. 2.2).  [^2] When $s$ is large, it diminishes influence of $u$ and reduces accuracy by increasing the variance of $\theta$. Since $u,s$ are intimately tied to the loss function $L,$ bounding $s$ is generally required per loss function and can depend on the model outputs. In this initial work, we consider $\ell^{2}$ loss functions ($\sum_{X,Y}(\hat{y}(x)-y)^{2}$) on classification tasks with model outputs and targets bounded by 1. Here we show these conditions give $s=1$.
Define the utility to be the negative loss so
$u=-L=\sum_{(x,y)\in(X,Y)}l(x,y,\theta)+r(\theta)$
where $l(x,y,\theta)$ is the loss from data point $(x,y)$, and $r(\theta)$ is a regularization function. For two neighboring datasets $(X,Y),(X^{\prime},Y^{\prime})$, let $(x,y)\in(X,Y)$ and $(x^{\prime},y^{\prime})\in(X^{\prime},Y^{\prime})$ denote the lone row of values that are not equal. We obtain $|u(X,Y,\theta)-u(X^{\prime},Y^{\prime},\theta)|=|l(x,y,\theta)-l(x^{\prime},y^{\prime},\theta)|$ (all terms of $L$ cancel except that of the differing data point); hence, bounding $s$ is achieved by Lipschitz or $\ell^{\infty}$ conditions on the per-data-point loss function $l$. Regularization $r$ does not affect sensitivity.
@@ 3 ExpM+NF Method
Corollary 3.2.
Sensitivity $s\leq 1$ for classification tasks using binary/one-hot encoded targets and logistic/soft-max outputs with $\ell^{2}$ loss and $u=-L.$
@@
Near optimal $\theta$ are found by sampling $\theta\sim p_{ExpM}(\theta)\propto\exp[u(X,Y,\theta)\varepsilon/2s]$ equipped with a pre-specified $\varepsilon$-DP guarantee. This step replaces an SGD scheme but fills the same role—it produces approximately optimal parameters for the ML model. Unfortunately, $p_{ExpM}$ is a historically intractable density for non-finite domains, so sampling directly is infeasible in most supervised learning applications. Our approach trains an auxiliary NF model (Section 2.2) to transform a Gaussian base density $p_{z}$ into our target density $p^{*}_{\theta}:=p_{ExpM}\propto\exp(u\varepsilon/(2s))$. The NF training (Section 2.2) is computationally well suited for approximating the ExpM density. The Reverse KL loss function (Eq. 1) only requires computing the determinant Jacobian (made tractable by the design of the NF) and $\log[p^{*}_{\theta}(\theta)Z]=u\varepsilon/(2s)$, essentially our loss function. Finally, supposing the NF output density is a good approximation of the ExpM density, a sample from the NF can be released with the $\varepsilon$-DP guarantee.
4 Examining ExpM+NF Potential on MIMIC-III
@@
The previous section outlines a new method for obtaining optimal parameters for an ML model, yet it is unclear if sampling parameters from an NF model can produce an accurate ML model especially, when ideally small $\varepsilon$ are used. To test the feasibility and show the promise of ExpM+NF method, we provide experimental results by training a logistic regression (LR) and a deep learning model (GRU-D) on MIMIC-III—a large, open, electronic health record database covering a diverse group of critical care patients at a medical center from 2001–2012. We use the binary prediction task benchmarks ICU Mortality and ICU Length of Stay ($>$ 3 days) from Wang et al. (Wang et al., 2020). Results of non-private LR and GRU-D (Che et al., 2018) models are provided by Wang et al. and Suriyakumar et al. (Suriyakumar et al., 2021) provide DPSGD (with RDP accountant) results on these tasks as well. To the best of our knowledge, Suriyakumar et al. (Suriyakumar et al., 2021) present the state of the art in differentially private ML on these MIMIC-III benchmarks. Unfortunately, their code is not provided.
We replicate these benchmarks as closely as possible. For all tasks and models (LR and GRUD) we implement a non-private baseline, DPSGD and ExpM+NF trained private versions. ExpM+NF is currently constrained to using $\ell^{2}$ loss (needed for the sensitivity bound). For the non-private and DPSGD models, we provide results for both binary cross entropy loss (BCE) and $\ell^{2}$ loss, as both are common for binary classification. For privacy methods, we vary $\varepsilon$ to see the accuracy-privacy tradeoff. Differentially private ML research generally considers $\varepsilon\in[1,10]$ (Ponomareva et al., 2023), hence we present plots of $0.5\leq\epsilon\leq 10$, and, investigate results for $\varepsilon$ as low as $1\mathrm{e}{-4}$. All experiments are coded in PyTorch.
@@
For each $\varepsilon$, we performed a randomized grid search of hyperparemeters fitting on the train set and evaluating on the dev (development, or validation set) set. For DPSGD, we use the PRV method of accounting via Opacus (Yousefpour et al., ; Yousefpour et al., 2021). Our NF models always use a Gaussian base distribution with $N(\vec{0},\text{diag}(\sigma^{2}))$, and we treat the constant $\sigma^{2}$ appearing on the diagonal of the covariance matrix as a hyperparameter. We use AUC as our accuracy metric as the task is binary classification. Starting conditions can lead to anomalously good/bad models for all methods (non-private, DPSGD, ExpM+NF) so we present the median across ten models evaluated on the hold out test set, each with different random seeds, to provide robustness to outliers.
@@
Further experimental set up details, including more information on the MIMIC-III dataset and benchmarks, how we performed the hyperparameter search, and how we present the results as the median across ten models, are in the Appendix C.
@@
Table 1: Median AUC results on MIMIC-III Mortality and Length of Stay prediction experiments
@@
4.1 Accuracy-Privacy Results
As our goal is to test if ExpM+NF is a viable training method, our results report accuracy (AUC) versus privacy ($\varepsilon$), assuming the sampling from the trained NF is the exponential mechanism. Under this assumption, our overall finding is that ExpM+NF substantially outperforms (our now state-of-the-art) DPSGD results in accuracy and privacy. Table 1 displays accuracy-versus-privacy results. In every experiment, ExpM+NF exhibits greater than 94% of the non-private models’ AUC for all $\varepsilon\geq 1\mathrm{e}-3$. ExpM+NF surpasses DPSGD in all experiments except $\varepsilon\geq 2$ with LR model on the Length of Stay task, in which all privacy methods achieve 99% of the non-private model’s AUC. Viewed through another lens, ExpM+NF maintains a greater AUC for $\varepsilon\geq 1\mathrm{e}-3$ than DPSGD for $\varepsilon=1,\delta=1\mathrm{e}-5$ in all MIMIC-III experiments, respectively. For both prediction tasks, the AUCs achieved by ExpM+NF of the GRU-D model are, for all $\varepsilon\geq 1\mathrm{e}-3$, better than the non-private LR or GRU-D models’ results reported by Suriyakumar (Suriyakumar et al., 2021). These results show that ExpM+NF is a viable training method even for $\varepsilon$ values orders of magnitude smaller (orders of magnitude better privacy) than is currently considered.
Appendix D provides many other findings including investigation of the NF distributions and ancillary results showing our DPSGD implementations push the previous state of the art for these MIMIC-III benchmarks.
4.2 Timing Results
Another aspect to the viability of this method is its impact on the training time performance (as it requires training an auxiliary NF model). Here we present timing results for ICU Mortality experiments in Table 2. Timing results for Length of Stay are similar and reside in Appendix D.3 along with more details of this timing experiment. All private training methods are slower than non-private (expected), and that when including the time to compute the noise_multiplier, DPSGD is slower than ExpM+NF for both LR and GRU-D training. The innovation here is that our ExpM+NF code computes (and tracks gradients through) the $\ell^{2}$ loss for the sampled parameters $\theta_{i}$ in parallel when computing the RKL loss for training the NF, rather than computing this sequentially. We used the best hyperparmeters for each model, $\varepsilon$, so batch size and epochs vary between the different implementations and privacy levels.
Table 2: ICU Mortality Mean Benchmark Timing Results
@@ 4.2 Timing Results
Training Loss Training Time Computing $\varepsilon$ Total Training Time Computing $\varepsilon$ Total

### v4

@@ 2.2. Normalizing Flows (NFs)
  Initialize $\beta_{\text{NF},0}$ randomly
  for $t$ in $[T]$Train Auxiliary NF Model} do
     Generate sample from base distribution
     $z_{i}\sim p_{z},i=1,...,N$
     Produce parameter sample & Jacobian
     $\theta_{i}=g_{\beta_{\text{NF},t}}(z_{i})$
@@
     Compute utility for each sample
@@
     Compute Reverse KL Loss
@@
     $\beta_{\text{NF},t+1}=\beta_{\text{NF},t}-\eta_{t}\nabla_{\beta_{\text{NF},t}}\mathcal{L}$
@@
  Output Model parameters, $\theta=g_{\beta_{\text{NF},T}}(z)$
Algorithm 1 Exp+NF
3. ExpM+NF
Figure 1 depicts the setting that we will consider, training and publicly releasing an ML algorithm on private data, and the proposed method, ExpM+NF, as described in Algorithm 1. For training data $(X,Y)$, we seek $\theta$ that minimizes a loss function $L(X,Y,\theta)$. ExpM requires specifying a utility function $u(X,Y,\theta)$ to be maximized in $\theta$; thus, $u=\phi\circ L$ for strictly decreasing $\phi$. Throughout, we will consider $u=-L$. To define the ExpM density, we must provide $s$, a bound on the sensitivity of $u$ (Defn. 2.2). (See Appendix C for more details on $\phi$ and $s$). In this initial work, we consider $\ell^{2}$ loss functions ($\sum_{X,Y}(\hat{y}(x)-y)^{2}$) on classification tasks with model outputs and targets bounded by 1.
@@ 3. ExpM+NF
Corollary 3.2.
$s\leq 1$ for classification tasks using binary/one–hot targets and logistic/soft–max outputs with $\ell^{2}$ loss and $u=-L.$
@@
Define the utility to be the negative loss so $u=-L=\sum_{(x,y)\in(X,Y)}l(x,y,\theta)+r(\theta)$ where $l(x,y,\theta)$ is the loss from data point $(x,y)$, and $r(\theta)$ is a regularization function. For two neighboring datasets $(X,Y),(X^{\prime},Y^{\prime})$, let $(x,y)\in(X,Y)$ and $(x^{\prime},y^{\prime})\in(X^{\prime},Y^{\prime})$ denote the lone row of values that are not equal. We obtain $|u(X,Y,\theta)-u(X^{\prime},Y^{\prime},\theta)|=|l(x,y,\theta)-l(x^{\prime},y^{\prime},\theta)|$ (all terms of $L$ cancel except that of the differing data point); hence, bounding $s$ is achieved by Lipschitz or $\ell^{\infty}$ conditions on the per–data–point loss function $l$. Regularization $r$ does not affect sensitivity.
@@
Near optimal $\theta$ are found by sampling
$\theta\sim p_{ExpM}(\theta)\propto\exp[u(X,Y,\theta)\varepsilon/2s]$
@@
equipped with a pre-specified $\varepsilon$-DP guarantee. This step replaces an SGD scheme but fills the same role, producing approximately optimal parameters. As sampling continuous $\theta\sim p_{ExpM}$ is intractable, we train an auxiliary NF model as discussed in Section 2.2. NF training is computationally well suited for approximating the ExpM density. The NF’s loss (Reverse KL ) only requires computing the determinant Jacobian (made tractable by the design of the NF) and $\log[p^{*}_{\theta}(\theta)Z]=u\varepsilon/(2s)$, essentially the target model’s loss function. Finally, if the NF output density is a good approximation of the ExpM density, a sample from the NF can be released with nearly $\varepsilon$-DP guarantee (see Theorem 7.3).
4. Testing ExpM+NF Model Training on MIMIC-III Data
@@
To test the feasibility of ExpM+NF, we provide experimental results by training a logistic regression (LR) and a deep recurrent network (GRU-D) on MIMIC-III—a large, open, electronic health record database. We use the binary prediction task benchmarks ICU Mortality and ICU Length of Stay ($>$ 3 days) from Wang et al. (wang2020mimic, ). Results of non-private LR and GRU-D (che2018recurrent, ) models are provided by Wang et al., and results from training via DPSGD (with RDP accountant) appear in Suriyakumar et al. (suriyakumar2021chasing, ). To the best of our knowledge, Suriyakumar et al. (suriyakumar2021chasing, ) present the state of the art in differentially private ML on these MIMIC-III benchmarks. Unfortunately, their code is not provided. We replicate these benchmarks as closely as possible.
For all tasks and models we implement a non-private (SGD) baseline, and DPSGD- and ExpM+NF-trained versions. ExpM+NF is currently constrained to using $\ell^{2}$ loss (needed for the sensitivity bound), while for SGD and DPSGD both $\ell^{2}$ and binary cross entropy (BCE) loss are used.
As $\varepsilon$ decreases, variance of the ExpM distribution increases; intuitively this should lead to a greater likelihood of sampling a model with poor accuracy. Differentially private ML research generally considers $\varepsilon\in[1,10]$ (ponomareva2023how, ). We present results for $\varepsilon\in[1\mathrm{e}{-4},10]$.
@@
For each $\varepsilon$, we performed a randomized grid search of hyperparemeters fitting on the train set and evaluating on the development set. For DPSGD, we use the PRV method of accounting via Opacus (pytorchOpacus, ; yousefpour2021opacus, ). Our NF models use base distribution $N(\vec{0},\text{diag}(\sigma^{2}))$, and we treat $\sigma$ as a hyperparameter. We use AUC as our accuracy metric. We present the median AUC on the held-out test set across ten models, each with different random seeds, to provide robustness to outliers. For ExpM+NF training, we also train ten NF models (each is a density over our target classifier). We use the median AUC of 1K samples per each of the ten NF models. Box plots and per-NF AUC densities are plotted in the Appendix Figure 7 for GRU-D ICU (and analogous plots for all experiments are in the code base, URL on front page).
As all our models are implemented in PyTorch, we use PyTorch’s benchmark suite which (only) reports the mean timing of $n$ runs, and only times the GPU training time (no data IO, etc.). We use $n=10$ runs (per task, model, training method, $\varepsilon$) on a single NVIDIA A100 80GB GPU and 128 CPUs with 2TB of shared RAM. For DPSGD and ExpM+NF, we report the mean across all $\varepsilon$. For the non-private GRU-D implementation, we report timing with batch normalization which was fastest.
@@
Further details, including information on the MIMIC-III data and benchmarks, our hyperparameter search, and how we present the results are in the Appendix D.
@@
Table 1. Median AUC results on MIMIC-III Mortality and Length of Stay prediction experiments
@@
4.1. Accuracy and Timing Results
The table of figures 1 displays accuracy as we vary $\varepsilon$. In every experiment, ExpM+NF exhibits greater than 94% of the non-private models’ AUC for all $\varepsilon\geq 1\mathrm{e}-3$. ExpM+NF surpasses DPSGD in all experiments except $\varepsilon\geq 2$ with the LR model on the Length of Stay task, in which all privacy methods achieve 99% of the non-private model’s AUC. Viewed through another lens, ExpM+NF maintains a greater AUC for $\varepsilon\geq 1\mathrm{e}-3$ than DPSGD for $\varepsilon=1,\delta=1\mathrm{e}-5$ in all MIMIC-III experiments, respectively. These results show that ExpM+NF is a viable training method even for $\varepsilon$ values orders of magnitude smaller than is currently considered. We do not claim ExpM+NF achieves $\varepsilon$-DP as this is not yet proven. Appendix E provides other findings including NF distributions and results showing our DPSGD implementations push the previous state of the art for these MIMIC-III benchmarks.
More discussion and takeaways of these results are provided in the Conclusion 8.
Timing results, which report only the computational time due to training, appear in Table 2 for ICU Mortality experiments. Timing results for Length of Stay are similar and reside in Appendix E.3 along with more details of this timing experiment. Non-private (SGD) is the fastest, as expected). When including computation of the noise_multiplier, DPSGD is slower than ExpM+NF for both LR and GRU-D training. The innovation here is, when computing the RKL loss for NF training, our ExpM+NF code computes (and tracks gradients through) the $\ell^{2}$ loss for all the sampled parameters $\theta_{i}=g(z_{i})$ in parallel, rather than computing this sequentially. We used the best hyperparmeters for each model, so batch size and epochs vary between the different tasks and $\varepsilon$ levels.
Table 2. ICU Mortality Mean Benchmark Timing Results
@@ 4.1. Accuracy and Timing Results
Training Loss Training Computing Total Training Computing Total
Time $\varepsilon$ Time $\varepsilon$